OpenStudy (anonymous):

5√3+√6/√6+√3? would the numerators cancel out leaving only the five? or is there more to this?

OpenStudy (saifoo.khan):

\[\Large \frac{5 \sqrt 3 + \sqrt6}{\sqrt6+\sqrt3} \to -3 +4\sqrt2\]

OpenStudy (anonymous):

hmm... could you explain in in further detail?

OpenStudy (saifoo.khan):

Can u solve now?

OpenStudy (anonymous):

ill try

OpenStudy (saifoo.khan):

wait i did a mistake,

OpenStudy (anonymous):

D:

OpenStudy (saifoo.khan):

\[\Large \frac{5 \sqrt 3 + \sqrt6}{\sqrt6+\sqrt3} \times \frac{\sqrt6-\sqrt3}{\sqrt6-\sqrt3}\]Now it's Accurate.

OpenStudy (anonymous):

i dont know where you got the second half of the problem... its only the first set of numbers

OpenStudy (saifoo.khan):

It's a method to solve these type of questions.

OpenStudy (anonymous):

oh neat

OpenStudy (saifoo.khan):

You have multiply and divide with the denominator but with different sign.

OpenStudy (anonymous):

im trying to get it worked out

OpenStudy (anonymous):

how do you multiply a √3 X -√3? thats what i dont understand

OpenStudy (saifoo.khan):

\[\sqrt 3 \times - \sqrt3 = -3 \]

OpenStudy (saifoo.khan):

The roots cancels out.

OpenStudy (anonymous):

ahhh. brilliant!

OpenStudy (saifoo.khan):

Thank you. heading to bed. C ya! :)

OpenStudy (anonymous):

thanks

OpenStudy (saifoo.khan):

please hit good answer/.

OpenStudy (anonymous):

sorry, wasnt watching